# DISTINQT: A Distributed Privacy Aware Learning Framework for QoS Prediction for Future Mobile and Wireless Networks

### v1

@@ Abstract
Beyond 5G and 6G networks are expected to support new and challenging use cases and applications that depend on a certain level of Quality of Service (QoS) to operate smoothly. Predicting the QoS in a timely manner is of high importance, especially for safety-critical applications as in the case of vehicular communications. Although until recent years the QoS prediction has been carried out by centralized Artificial Intelligence (AI) solutions, a number of privacy, computational, and operational concerns have emerged. Alternative solutions have been surfaced (e.g. Split Learning, Federated Learning), distributing AI tasks of reduced complexity across nodes, while preserving the privacy of the data. However, new challenges rise when it comes to scalable distributed learning approaches, taking into account the heterogeneous nature of future wireless networks. The current work proposes DISTINQT, a privacy-aware distributed learning framework for QoS prediction. Our framework supports multiple heterogeneous nodes, in terms of data types and model architectures, by sharing computations across them. This, enables the incorporation of diverse knowledge into a sole learning process that will enhance the robustness and generalization capabilities of the final QoS prediction model. DISTINQT also contributes to data privacy preservation by encoding any raw input data into a non-linear latent representation before any transmission. Evaluation results showcase that our framework achieves a statistically identical performance compared to its centralized version and an average performance improvement of up to 65% against six state-of-the-art centralized baseline solutions in the Tele-Operated Driving use case.
@@
Beyond 5G, 6G, QoS Prediction, Tele-Operated Driving, Distributed Learning
@@ I Introduction
Motivated by these open challenges, the current work proposes DISTINQT, a novel privacy-aware distributed learning framework based on sequence-to-sequence autoencoders [21], capable of providing accurate predictions to end users for a future time horizon. More specifically, the DISTINQT’s architecture extends upon our previous work [22], preserving data privacy by distributing (among distinct nodes) a number of NN-Encoders, capable of encoding input data into a non-linear latent representation before any transmission. DISTINQT also allows different architectural options for each NN-Encoder, depending on the data types collected at the different nodes. This flexibility enables the incorporation of diverse knowledge and model architectures into a sole learning process that will enhance the robustness and generalization capabilities of the final QoS prediction model.
@@
  • We propose a distributed learning framework for QoS prediction, sharing computations of lower complexity among different nodes, training a complex NN model.
@@
  • Our framework is privacy-aware, contributing to data privacy preservation by encoding raw input data into a non-linear latent representation before any transmission.
@@ II Related Work
There is also a number of works related to the application of distributed learning in wireless networks for a variety of use cases including computational offloading, model partitioning and energy efficiency. In works [29] and [30] the authors propose a fine-grained adaptive partitioning scheme to divide a source DNN in partitions. Two integrating options are proposed including a local execution at an end device, or partition offloading to one or multiple powerful nodes. Another work [31] investigates DNN inference offloading in a 5G environment, aiming to increase the prediction request acceptance or to minimize the energy consumption of edge devices. Zeng et. al [32], propose CoEdge, a distributed DNN computing system that orchestrates cooperative prediction over heterogeneous edge devices towards minimizing the system energy consumption while promising response latency requirement. In another work [33] the authors propose DeCNN, a prediction approach that uses decoupled Convolutional NN (CNN) structure to optimize multi-level parallelism for distributed prediction on end devices. Additionally in [34], the authors propose DeepHome, a framework that can distribute NN inference tasks to multiple heterogeneous devices inside a home.
Finally, only a limited number of works focus on the application of distributed learning for QoS prediction in mobile networks. Zhang et.al [18] propose a privacy-preserving QoS prediction, leveraging federated learning techniques, along with several efficiency improvements to reduce system overhead. The authors in [19] propose MultiFed, which is a multi-centre FL framework for QoS prediction via a cloud–edge collaboration mechanism. The work in [20] discusses a personalized federated tensor factorization framework for distributed privacy-preserving IoT services QoS prediction. However, these works assume identical data types and NN architectures across nodes, while considering differences only in data distribution and size.
As already stated in the Introduction, there is a gap in the state-of-the-art works when it comes to distributed learning approaches for QoS prediction taking into account both common and heterogeneous nodes, in terms of data types and model architectures, along with data privacy and knowledge sharing. The current work extends [22], by enabling the distribution of the learning apart from the inference/prediction process. DISTINQT is the first one that proposes a distributed learning framework for QoS prediction, sharing computations of lower complexity among different nodes across the network, supporting network entities with similar and heterogeneous data types and model architectures. Our framework contributes to data privacy preservation by encoding raw input data into a non-linear latent representation before any transmission.
@@ III DISTINQT Framework
This section provides an extensive description of the DISTINQT framework towards a privacy preserving distributed learning scheme for QoS prediction. DISTINQT could be effectively applied to any wireless network environment comprised of distinct types of Network Entities (NETs), such as UE, BS, MEC, Cloud etc. DISTINQT follows an hierarchical structure, enabling distibuted learning in large-scale network environments. Let $E$ denote the number of distinct NETs. Each NET $e\in\{1,...,E\}$ may be comprised of multiple workers. All workers of a specific NET share the same input features (data types) but the number and distribution of data samples may differ. However, workers of different NETs collect and process different input features related with a learning task.
@@ III-A Role Assignment
The DISTINQT framework is realized in a synchronized manner and is capable of providing QoS predictions for a specific time horizon. At the beginning of the learning phase, each NET is assigned with one of the following roles:
@@
Local NET Aggregator: A local aggregator is assigned to each NET, acting as an intermediate point in the distributed forward and backpropagation processes, by collecting and transmitting all the necessary inputs, aggregating and updating the trainable weights of the involved workers of a specific NET.
@@ III-B Neural Network Architecture and Distribution
Fig. 1 depicts a high level overview of the DISTINQT’s NN architecture that will be distributed among all involved workers of different NETs. The DISTINQT framework is based on sequence-to-sequence Autoencoder NN architectures [21, 35, 36, 37], designed to encode an input sequence into a compressed representation. Apart from the inherent dimensionality reduction, these architectures also contribute to privacy preservation, by encoding the initial raw data before any transmission.
@@
NN-Encoder: Each worker is equipped with one NN-Encoder. Let $\mathsf{K}_{e}$ denote the total number of workers of NET $e$, $N_{enc}=\sum_{e=1}^{E}\mathsf{K}_{e}$ denote the total number of NN-Encoders in the system. Each NN-Encoder may be comprised of multiple encoding layers. An NN-Encoder processes an input sequence and encodes it into a fixed-length vector, named as context vector, while applying a non-linear activation function (e.g. the rectified linear activation function - ReLU). This preserves data privacy, since the mapping of the input sequence to a non-linear latent representation is of higher complexity and as such cannot be easily reversed by any other worker or snooper [38, 39]. The workers of the same NET employ identical NN-Encoders, in terms of number and type of encoding layers, as well as number of neurons per layer. Indicative examples of such NN-Encoder types include: Long-Short Term Memory, Transformer, Convolutional encoders etc. The NN-Encoders used by distinct NETs may differ in their NN structure, depending on the use case, data types (e.g. visual, aural, numerical), and computational capabilities of the workers. DISTINQT’s flexibility in architecture configuration could enable and support a multi-modal learning process [40].
@@
Merging Layer: This layer is responsible for applying a merging function (e.g. summation, concatenation, average) to the context vectors produced by all workers, including the coordinator. The merging is performed between interconnected (associated) workers resided on different NETs (more details in Section III-C).
NN-Decoder: The NN-Decoder is responsible for constructing the future sequence over the selected time horizon, based on the output of the Merging Layer. It should be noted that the NN-Decoder does not attempt to reconstruct any initial input sequence.
@@
III-C Methodology
@@
Step \⃝raisebox{-0.9pt}{1}: Each worker periodically collects a sequence of input features. The logging frequency of the input features $\tau_{e}$ ($e\in\{1,...,E\}$) may vary per NET and it is expressed through a discrete time unit, the timestep. In each batch loop $b$, all NN-Encoders process a batch $B$ of input sequences. The batch size $s$ is selected by NET $c$, based on the prediction task, and it is the same for all workers. Each input sequence initiates a forward pass process. This sequence includes historical features for $h_{e}$ consecutive past timesteps, comprising a historical time horizon with duration $H_{e}$ ($H_{e}=h_{e}\cdot\tau_{e}$).
@@
Step \⃝raisebox{-0.9pt}{2}: This step includes the merging of the context vectors that is performed by the coordinator NET, which keeps track of the interconnections. The merging is realized only among the interconnected workers of different NETs. As an example, consider a wireless mobile network of three NETs, namely a UE, a BS, and a MEC, with three, two and one workers of each type, shown in Fig. 3. The different line colors indicate the different interconnections (three in total). The MEC worker acts as the coordinator. Fig. 4 maps this wireless network to DISTINQT and illustrates the merging process. The context vector of each UE worker will be merged with the context vector of the connected (associated) BS worker and consequently with the context vector of the associated (to the respective BS) MEC-worker. As a result three merged vectors will be produced by applying a merging function to the following vectors: [$\textbf{v}_{t}^{(\textrm{UE},1)}$, $\textbf{v}_{t}^{(\textrm{BS},1)}$, $\textbf{v}_{t}^{(\textrm{MEC},1)}$], [$\textbf{v}_{t}^{(\textrm{UE},2)}$, $\textbf{v}_{t}^{(\textrm{BS},1)}$, $\textbf{v}_{t}^{(\textrm{MEC},1)}$], [$\textbf{v}_{t}^{(\textrm{UE},3)}$, $\textbf{v}_{t}^{(\textrm{BS},2)}$, $\textbf{v}_{t}^{(\textrm{MEC},1)}$].
@@
Let $A_{t}$ denote the total number of merged vectors produced by the merging layer at timestep $t$. All merged vectors are of the same length, denoted by $M$. Let $m_{t,i}^{(a)}$ with $a\in\{1,...,A_{t}\}$ and $i\in\{1,...,M\}$, denote the $i^{th}$ element of the merged vector of interconnection $a$ at timestep $t$. Consequently, the output of the merging layer, associated with interconnection $a$, is the vector $\textbf{m}^{(a)}_{t}=[m_{t,1}^{(a)},...,m_{t,M}^{(a)}]^{T}$. In the example given by Fig. 4, the produced merged vectors are the $\textbf{m}^{(1)}_{t}$, $\textbf{m}^{(2)}_{t}$ and $\textbf{m}^{(3)}_{t}$.
Steps \⃝raisebox{-0.9pt}{3}-\⃝raisebox{-0.9pt}{5}: For each interconnection $a$, the merged vector $\textbf{m}^{(t)}_{a}$ is provided as input to the NN-Decoder. As mentioned, the NN-Decoder constructs the future sequence of the selected prediction time horizon $\mathsf{P}$ ($\mathsf{P}=p\cdot\uptau_{c}$) and it is comprised of $p$ prediction horizon timesteps in total. This prediction horizon timestep is selected by the coordinator and is denoted by $\uptau_{c}$.
$E$
Number of distinct NETs
$\mathsf{K}_{e}$
Number of workers of NET $e$
$B$
Batch of input sequences
$b$
Batch loop
$s$
Batch size
$\tau_{e}$
Logging frequency timestep of NET $e$
$\uptau_{c}$
Prediction horizon timestep selected by NET $c$
$t$
Timestep of the current forward pass
$h_{e}$
Number of timesteps in a historical time horizon $H_{e}$
$p$
Number of timesteps in a prediction time horizon $\mathsf{P}$
$H_{e}$
Duration of historical time horizon of NET $e$
$\mathsf{P}$
Duration of the prediction time horizon
$N_{enc}$
Total number of Encoding layers (NN-Encoders)
$N_{e}$
Length of the context vector of NET $e$
$\nu^{(e,k)}_{t,i}$
The $i^{th}$ element of the context vector of worker $k$ of NET $e$ at timestep $t$
$\textbf{v}^{(e,k)}_{t}$
Context vector of worker $k$ of NET $e$ at timestep $t$
$\textbf{v}^{(c)}_{t}$
Context vector of coordinator $c$ at timestep $t$
$A_{t}$
Total number of merged vectors at timestep $t$
$M$
Length of the merged vector
$m_{t,i}^{(a)}$
The $i^{th}$ element of the merged vector of interconnection $a$ at timestep $t$
$\textbf{m}^{(a)}_{t}$
Vector output of the merging layer
$\hat{\textbf{y}}^{(a)}_{t}$
Predicted output sequence of interconnection $a$ at timestep $t$
$\hat{\textbf{y}}_{b}$
Set of predicted output sequences at batch loop $b$
$\textbf{w}_{b}^{(e,k)}$
The weight updates of worker $k$ of NET $e$ at batch loop $b$
$\textbf{w}_{b}^{(e)}$
Aggregated trainable weights of NET $e$ at batch loop $b$
@@
Step \⃝raisebox{-0.9pt}{6}: Once the forward pass processes of batch loop $b$ are complete, the predicted output sequence $\hat{\textbf{y}}_{b}$ along with the groundtruth sequence will be used, in order to minimize the loss function of the DISTINQT framework and perform the backpropagation process. During the backpropagation, the gradients of each layer (partial derivatives of the loss function w.r.t the trainable parameters) are computed by recursively applying the chain rule up to the NN-Encoder of worker $c$. The computed gradients will be used in order to update the trainable weights of worker $c$.
Step \⃝raisebox{-0.9pt}{7}: The computed gradients of the NN-decoder are forwarded to the Local NET Aggregators of the involved NETs to continue a distributed backpropagation process. The Local NET Aggregators broadcast the received gradients to the workers of the associated NET to update their trainable weights using their own optimizers. Let $\textbf{w}_{b}^{(e,k)}$ denote the weight updates of worker $k$ of NET $e$ at the end of batch loop $b$.
Step \⃝raisebox{-0.9pt}{8}: All workers send their updated weights to the respective Local NET Aggregator for aggregation based on an averaging function. The Local NET Aggregator of each NET $e$ updates the Global NN-Encoder with the aggregated weights $\textbf{w}_{b}^{(e)}$, given by:
$\textbf{w}_{b}^{(e)}=\frac{\sum_{k=1}^{\mathsf{K}_{e}}\textbf{w}_{b}^{(e,k)}}{\mathsf{K}_{e}}\vspace{-8pt}$ (1)
Step \⃝raisebox{-0.9pt}{9}: The Local NET Aggregator broadcasts the aggregated trainable weights $\textbf{w}_{b}^{(e)}$ back to the respective workers to update their NN-Encoder. This step completes a batch loop of the learning process. It should be noted that after this step, all workers of a NET will have the same NN-Encoder, i.e. same trainable weights.
Steps \⃝raisebox{-0.9pt}{1}-\⃝raisebox{-0.9pt}{9} are repeated for each batch loop and for each epoch of the learning process until convergence.
Finally, it should be mentioned that since the learning process is repeated for each batch loop, it could potentially lead to an increase in the communication overhead. This overhead could be potentially mitigated by applying effective compression techniques for NNs, such as connection pruning, weight quantization, knowledge distillation [42]. However, since this aspect is not covered in this work, we assume that there are no strict limitations in communication resources.

### v2

@@ Abstract
Beyond 5G and 6G networks are expected to support new and challenging use cases and applications that depend on a certain level of Quality of Service (QoS) to operate smoothly. Predicting the QoS in a timely manner is of high importance, especially for safety-critical applications as in the case of vehicular communications. Although until recent years the QoS prediction has been carried out by centralized Artificial Intelligence (AI) solutions, a number of privacy, computational, and operational concerns have emerged. Alternative solutions have surfaced (e.g. Split Learning, Federated Learning), distributing AI tasks of reduced complexity across nodes, while preserving the privacy of the data. However, new challenges rise when it comes to scalable distributed learning approaches, taking into account the heterogeneous nature of future wireless networks. The current work proposes DISTINQT, a novel multi-headed input privacy-aware distributed learning framework for QoS prediction. Our framework supports multiple heterogeneous nodes, in terms of data types and model architectures, by sharing computations across them. This enables the incorporation of diverse knowledge into a sole learning process that will enhance the robustness and generalization capabilities of the final QoS prediction model. DISTINQT also contributes to data privacy preservation by encoding any raw input data into highly complex, compressed, and irreversible latent representations before any transmission. Evaluation results showcase that DISTINQT achieves a statistically identical performance compared to its centralized version, while also proving the validity of the privacy preserving claims. DISTINQT manages to achieve a reduction in prediction error of up to 65% on average against six state-of-the-art centralized baseline solutions presented in the Tele-Operated Driving use case.
@@
Beyond 5G, 6G, QoS Prediction, Tele-Operated Driving, Privacy, Distributed Learning
@@ I Introduction
Motivated by these open challenges, the current work proposes DISTINQT, a novel multi-headed input [21] privacy-aware distributed learning framework based on sequence-to-sequence autoencoders [22], capable of providing accurate predictions to end users for a future time horizon. More specifically, the DISTINQT’s architecture extends upon our previous work [23], preserving data privacy by distributing (among distinct nodes) a number of NN-Encoders, capable of encoding input data into highly complex, compressed, and irreversible latent representations before any transmission. DISTINQT also allows different architectural options for each NN-Encoder, depending on the data types collected at the different nodes. This flexibility enables the incorporation of diverse knowledge and model architectures into a sole learning process that will enhance the robustness and generalization capabilities of the final QoS prediction model.
@@
  • We propose a novel distributed learning framework for QoS prediction, sharing computations of lower complexity among different nodes, training a complex multi-headed input NN architecture.
@@
  • Our framework is privacy-aware, contributing to data privacy preservation by encoding raw input data into hlhighly complex, compressed, and irreversible latent representations before any transmission.
  • Our framework is evaluated in the Tele-Operated Driving use case achieving a statistically identical performance compared to its centralized counterpart and an average reduction in prediction error up to 65% against six state-of-the-art centralized baseline solutions.
@@ II Related Work
There is also a number of works related to the application of distributed learning in wireless networks for a variety of use cases including computational offloading, model partitioning and energy efficiency. In works [30] and [31] the authors propose a fine-grained adaptive partitioning scheme to divide a source DNN in partitions. Two integrating options are proposed including a local execution at an end device, or partition offloading to one or multiple powerful nodes. Another work [32] investigates DNN inference offloading in a 5G environment, aiming to increase the prediction request acceptance or to minimize the energy consumption of edge devices. Zeng et. al [33], propose CoEdge, a distributed DNN computing system that orchestrates cooperative prediction over heterogeneous edge devices towards minimizing the system energy consumption while promising response latency requirement. In another work [34] the authors propose decoupled Convolutional NN (DeCNN) structure to optimize multi-level parallelism for distributed prediction on end devices. Additionally in [35], the authors propose DeepHome, a framework that can distribute NN inference tasks to multiple heterogeneous devices inside a home.
Finally, only a limited number of works focus on the application of distributed learning for QoS prediction in mobile networks. Zhang et.al [18] propose a privacy-preserving QoS prediction, leveraging federated learning techniques, along with several efficiency improvements to reduce system overhead. The authors in [19] propose MultiFed, which is a multi-centre FL framework for QoS prediction via a cloud–edge collaboration mechanism. The work in [20] discusses a personalized federated tensor factorization framework for distributed privacy-preserving Internet of Things (IoT) services QoS prediction. However, these works assume identical data types and NN architectures across nodes, while considering differences only in data distribution and size.
As already stated in the Introduction, there is a gap in the state-of-the-art works when it comes to distributed learning approaches for QoS prediction taking into account both common and heterogeneous nodes, in terms of data types and model architectures, along with data privacy and knowledge sharing. The current work extends [23], by enabling the distribution of the learning apart from the inference/prediction process. DISTINQT is the first one that proposes a distributed learning framework for QoS prediction, sharing computations of lower complexity among different nodes across the network, supporting network entities with similar and heterogeneous data types and model architectures. Our framework contributes to data privacy preservation by encoding raw input data into hlhighly complex, compressed, and irreversible latent representations before any transmission.
@@ III DISTINQT Framework
This section provides an extensive description of the DISTINQT framework towards a privacy preserving distributed learning scheme for QoS prediction. DISTINQT could be effectively applied to any wireless network environment comprised of distinct types of Network Entities (NETs), such as User Equipment (UE), Base Station (BS), Multi-access Edge Computing (MEC), Cloud etc. DISTINQT follows an hierarchical structure, enabling distributed learning in large-scale network environments. Let $E$ denote the number of distinct NETs. Each NET $e\in\{1,...,E\}$ may be comprised of multiple workers. All workers of a specific NET share the same input features (data types) but the number and distribution of data samples may differ. However, workers of different NETs collect and process different input features related with a learning task.
@@ III-A Role Assignment
The DISTINQT framework is realized in a synchronized manner and is capable of providing QoS predictions for a specific time horizon. At the beginning of the learning phase, each NET is assigned with one of the following roles based on data samples availability:
@@
Local NET Aggregator: A local aggregator is assigned to each NET by the Coordinator NET, acting as an intermediate point in the distributed forward and backpropagation processes, by collecting and transmitting all the necessary inputs, aggregating and updating the trainable weights of the involved workers of a specific NET.
@@ III-B Neural Network Architecture and Distribution
Fig. 1 depicts a high level overview of the DISTINQT’s NN architecture that will be distributed among all involved workers of different NETs. DISTINQT, is a novel framework that introduces a multi-headed input [21] NN architecture based on sequence-to-sequence Autoencoders [22, 36, 37, 38]. DISTINQT is designed to encode multiple heterogeneous raw input sequences in parallel into highly complex, compressed, and irreversible latent representations (details in Section III-C). As a result, apart from the inherent dimensionality reduction, DISTINQT also contributes to privacy preservation, by encoding the initial raw data before any transmission.
@@
NN-Encoder: Each worker is equipped with one NN-Encoder. Let $\mathsf{K}_{e}$ denote the total number of workers of NET $e$, $N_{enc}=\sum_{e=1}^{E}\mathsf{K}_{e}$ denote the total number of NN-Encoders in the system. Each NN-Encoder may be comprised of multiple encoding layers. An NN-Encoder processes an input sequence and encodes it into a fixed-length vector, named as context vector. The workers of the same NET employ identical NN-Encoders, in terms of number and type of encoding layers, as well as number of neurons per layer. Indicative examples of such NN-Encoder types include: Long-Short Term Memory, Transformer, Convolutional encoders etc. DISTINQT shares computations of lower complexity, as the NN-Encoders used by distinct NETs may differ in their NN structure, depending on the use case, data types (e.g. visual, aural, numerical), and computational capabilities of the workers. DISTINQT’s flexibility in architecture configuration could enable and support a multi-modal learning process [39].
@@
Merging Layer: This layer is responsible for applying a merging function (e.g. summation, concatenation, average) to the context vectors produced by all workers, including the coordinator. The merging is performed between interconnected (associated) workers resided on different NETs (more details in Section III-D).
NN-Decoder: The NN-Decoder is responsible for constructing the future sequence over the selected time horizon, based on the output of the Merging Layer. Contrary to its name, it should be highly noted that the NN-Decoder does not reconstruct any initial input sequence and as such cannot be considered as a privacy leak.
@@
III-C Data Privacy Preservation
In addition to distributing a complex learning process among the involved workers of different NETs, the DISTINQT framework also aims to preserve the privacy of the initial raw data. DISTINQT allows solely the exchange of trainable parameters, gradients and context vectors between the participating NETs (Section III-D). Similarly to FL, which was designed with privacy-awareness as a core principle, trainable parameters and gradients exchanges are not perceived as privacy violations. In the case of context vectors, a first level of de facto privacy is maintained, as they are complex latent representations of the initial raw input sequences. On top of that, each NN-Encoder of the DISTINQT further enhances the complexity of the produced context vectors by introducing non-linearity through the application of an appropriate activation function (e.g. the rectified linear activation function - ReLU) [40, 41].
To further enhance the privacy of the raw data, two other methods are considered, including the application of regularization methods to the trainable parameters of each NN-Encoder and the raw input data normalization. Regularizers highly contribute to data privacy preservation by ensuring that the NN-Encoders will focus on capturing underlying data patterns, rather than memorizing specific details of the initial raw input sequences [42]. Normalization indirectly contributes to data privacy preservation by not exposing the value ranges of the raw input data [43]. Overall, the application of all the aforementioned methods to produce the context vectors preserve the privacy of the raw input data sequences. Consequently, the produced context vectors cannot be reversed by any other worker or snooper that has gained access to the respective NN-Encoders (Section V-C).
III-D Methodology
@@
[…]
@@
[…]
@@
Let $A_{t}$ denote the total number of merged vectors produced by the merging layer at timestep $t$. All merged vectors are of the same length, denoted by $M$. Let $m_{t,i}^{(a)}$ with $a\in\{1,...,A_{t}\}$ and $i\in\{1,...,M\}$ denote the $i^{th}$ element of the merged vector of interconnection $a$ at timestep $t$. Consequently, the output of the merging layer, associated with interconnection $a$, is the vector $\textbf{m}^{(a)}_{t}=[m_{t,1}^{(a)},...,m_{t,M}^{(a)}]^{T}$. In the example given by Fig. 4, the produced merged vectors are the $\textbf{m}^{(1)}_{t}$, $\textbf{m}^{(2)}_{t}$ and $\textbf{m}^{(3)}_{t}$.
[…]
$E$ Number of distinct NETs
$\mathsf{K}_{e}$ Number of workers of NET $e$
$B$ Batch of input sequences
$b$ Batch loop
$s$ Batch size
$\tau_{e}$ Logging frequency timestep of NET $e$
$\uptau_{c}$ Prediction horizon timestep selected by NET $c$
$t$ Timestep of the current forward pass
$h_{e}$ Number of timesteps in a historical time horizon $H_{e}$
$p$ Number of timesteps in a prediction time horizon $\mathsf{P}$
$H_{e}$ Duration of historical time horizon of NET $e$
$\mathsf{P}$ Duration of the prediction time horizon
$N_{enc}$ Total number of Encoding layers (NN-Encoders)
$N_{e}$ Length of the context vector of NET $e$
$\nu^{(e,k)}_{t,i}$ The $i^{th}$ element of the context vector of worker $k$ of NET $e$ at timestep $t$
$\textbf{v}^{(e,k)}_{t}$ Context vector of worker $k$ of NET $e$ at timestep $t$
$\textbf{v}^{(c)}_{t}$ Context vector of coordinator $c$ at timestep $t$
$A_{t}$ Total number of merged vectors at timestep $t$
$M$ Length of the merged vector
$m_{t,i}^{(a)}$ The $i^{th}$ element of the merged vector of interconnection $a$ at timestep $t$
$\textbf{m}^{(a)}_{t}$ Vector output of the merging layer
$\hat{\textbf{y}}^{(a)}_{t}$ Predicted output sequence of interconnection $a$ at timestep $t$
$\hat{\textbf{y}}_{b}$ Set of predicted output sequences at batch loop $b$
$\textbf{w}_{b}^{(e,k)}$ The weight updates of worker $k$ of NET $e$ at batch loop $b$
$\textbf{w}_{b}^{(e)}$ Aggregated trainable weights of NET $e$ at batch loop $b$
@@
[…]
$\textbf{w}_{b}^{(e)}=\frac{\sum_{k=1}^{\mathsf{K}_{e}}\textbf{w}_{b}^{(e,k)}}{\mathsf{K}_{e}}.\vspace{-8pt}$ (1)
[…]